A rectangle with it dimensions in centimetres Length up(3y+2)cm and down(x+7)cm and breadth left 2xcm and right (x+y+1)cm
(a) find the value of x and y
(b) the dimension of the rectangle

Found 3 solutions by josgarithmetic, MathTherapy, ikleyn:
Answer by josgarithmetic(39620)

Straighten the description. A rectangle has two dimensions and you named numbers for four quantities. Tell how your "left" and "right" numbers fit into this rectangle.

Answer by MathTherapy(10552)


A rectangle with it dimensions in centimetres Length up(3y+2)cm and down(x+7)cm and breadth left 2xcm and right (x+y+1)cm
(a) find the value of x and y
(b) the dimension of the rectangle
 


Answer by ikleyn(52798)

.
A rectangle with it dimensions in centimeters Length up (3y+2) cm and down (x+7) cm and breadth left 2x cm and right (x+y+1) cm.
(a) find the value of x and y
(b) the dimension of the rectangle
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

A rectangle has the opposite sides of equal length.
It gives you the system of two equations 

3y + 2 = x + 7,
2x = x + y + 1.

Solve it by any method you know to get the answer provided by the tutor "MathTherapy"

The writing by "josgarithmetic" is irrelevant. Ignore it.